OpenStudy (anonymous):

here's a fun question I got a while ago: What is the smallest number "n" such that: n /2 is a perfect square n/3 is a perfect cube and n/5 is a perfect fifth power?

OpenStudy (anonymous):

4/2=2 9/3=3 25/5=5

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

72

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

a multiple of it but

OpenStudy (anonymous):

a multiple that is diviisable by 5 I think

OpenStudy (anonymous):

but im going to sleep , tired

OpenStudy (anonymous):

are pagal ho gya hai kya

OpenStudy (radar):

O.K. Joe, don't keep us in suspense here too long!

OpenStudy (anonymous):

I found it

OpenStudy (anonymous):

2^15*3^10*5^36

OpenStudy (anonymous):

28156757354736328125000000000000000

OpenStudy (anonymous):

Sadly, that's the smallest n

OpenStudy (anonymous):

The smallest number is 0.

OpenStudy (anonymous):

Bah, i should have said non-zero n, good job abtrehearn for catching that mistake Heromiles is the closest, he has the right idea, but the exponent for 5 can be smaller and still work. Here is a pdf of my solution

OpenStudy (radar):

Thanks joemath(pi) now we can all get back to work lol.